Hubert is three years older than Jovie. The sum of their ages is 51. Find the age of Jovie
Answer by macston(5194)

H=Hubert's age; J=Jovie's age
H=J+3 yrs
H+J=51 yrs Substitute for H.
J+3yrs+J=51yrs Subtract 3 yrs from each side.
2J=48yrs Divide each side by 2
J=24 yrs ANSWER: Jovie is 24 yrs old.
CHECK:
H=J+3yrs=24yrs+3yrs=27yrs
H+J=51yrs
27yrs+24Yrs=51yrs
51yrs=51yrs